# Proto-LeakNet: Towards Signal-Leak Aware Attribution in Synthetic Human Face Imagery

### v1

@@ 3.5 Evaluation Metrics
    $\text{EER}=\min_{\tau}\big|\text{FAR}(\tau)-\text{FRR}(\tau)\big|,$ (19)
    where $\tau$ is the decision threshold. Lower EER indicates better separation between known and unknown samples.
@@
4 Dataset
All experiments are conducted on the WILD dataset [bongini2025wild], a large-scale benchmark for synthetic image attribution explicitly designed for closed- and open-set evaluation. It contains 20,00020{,}000 high-quality head-and-shoulder portraits generated by twenty distinct text-to-image models, including diffusion- and GAN-based systems such as Stable Diffusion, Midjourney, DALL-E, and StyleGAN. No image, prompt, or generator overlaps across splits, ensuring complete separation and preventing any data leakage. This dataset was selected for its diversity of generated samples, its unbiased distribution across generative models, and its inclusion of a rigorously designed post-processed test set for robustness evaluation.
4.1 Closed-Set
The closed set comprises 10,00010{,}000 images from ten generators, each contributing 1,0001{,}000 samples created from unique randomized prompts controlling facial and contextual attributes. Data are split into 5,0005{,}000 training, 2,0002{,}000 validation, and 3,0003{,}000 test samples, partitioned strictly at the prompt level to avoid cross-split repetition. Examples of raw images are shown in Figure 4.
4.2 Post-Processed Closed-Set
To evaluate robustness, half of the closed-set samples undergo up to three random transformations (compression, geometric, or photometric adjustments). Parameters are stochastically sampled to prevent duplicate augmentations. These post-processed images are used exclusively at inference, providing a controlled robustness benchmark against real-world degradation. An example of these post-processed samples can be visualized in Figure 4.
4.3 Open-Set
The open set includes 10,00010{,}000 additional portraits generated by ten unseen models, fully disjoint from those used for training. This split evaluates the model’s ability to generalize across architectures and prompt domains, measuring attribution consistency on previously unseen generators.
Figure 4: Samples from the WILD dataset, for each image in the first row, its column represent its various steps of post processing.
@@ 5 Experimental Results
This section presents the experiments conducted to evaluate the effectiveness of the proposed Proto-LeakNet framework in Section 3 across different scenarios. We compare our method against state-of-the-art approaches introduced in Section 2 using the same dataset and experimental conditions. Finally, we include ablation studies to assess the contribution of each component in our pipeline.
@@ 5.1 Results on Closed Set
We first evaluated Proto-LeakNet under the closed-set configuration to verify its ability to learn and capture generator-specific signal-leak patterns. This experiment establishes whether the model forms compact, well-separated latent clusters when all generator classes are known during training. Among all other methods we additionally reported the best three models reported on the WILD dataset [bongini2025wild] all trained in our exact conditions. As shown in the first block of Table 1, Proto-LeakNet achieves the highest Macro AUC among all compared methods on the raw closed-set, confirming its superior latent discriminability. In the first block of Table 1 we can closely inspect the single AUCs for each class to observe how our model model achieves superior performances overall. The combination of temporal attention and prototype supervision proves essential for this structured separation. Attention emphasizes the most informative diffusion timesteps, while prototypes enforce geometric consistency by pulling samples toward class centroids and repelling those of different generators. To further support these findings, Figure 5 reports the distributions of the Top-1 Accuracy on the raw Closed Set showing that our models still achieves a competitive accuracy compared to other methods. Together, these mechanisms yield a compact and interpretable latent manifold, as illustrated in Figure 2, demonstrating that Proto-LeakNet effectively captures intrinsic signal-leak biases that underpin robust generator attribution.
@@ 5.2 Results on Post-Processed Closed Set
To evaluate robustness against real-world degradations, we tested all models on progressively post-processed closed-set samples, applying up to three levels of perturbations described in Section 4. The experiment used pretrained checkpoints from the models trained on raw closed-set configuration described before. During this experiment, the models are trained exclusively on the raw closed-set samples, while post-processed images from Steps 1 to 3 are introduced only at inference time, without any additional fine-tuning. As shown in the second, third and fourth block in Table 1, all models performance decay due to the higher number of perturbance introduced but Proto-LeakNet consistently maintains the highest Macro AUC across all perturbation levels, confirming its resilience to visual degradation. This robustness arises from the model’s reliance on latent-domain signal-leak cues rather than pixel-level textures. Since signal-leak originates from generator-specific residuals embedded in the latent representation, these cues persist even after post-processing operations that distort pixel information. In contrast, competing methods operating in the image domain exhibit progressive performance collapse as perturbations increase, indicating their dependence on surface-level artifacts. Table 1 allows us to closely inspect single AUCs per class to assess proper attribution and in all three steps our proposed method achieves best performances. To support the results obtained in this experiment, Figure 5 reports the distribution of Top-1 Accuracy throughout all post processing steps, where Proto-LeakNet visibly achieves the highest accuracy proving to be the most robust to post-process attacks. Overall, these results demonstrate that Proto-LeakNet effectively captures stable generator-specific biases at the latent level, preserving attribution accuracy even under aggressive degradations that typically erode pixel-based forensic evidence.
Figure 5: Histogram of the Top-1 accuracies distributions of each method per class portraying how each class accuracy behaves with more degrees of degradation to the images.
Table 1: AUC results (%) for individual class per method including Macro AUC starting from results on raw closed-set until individual results for each post-processed experiment from step 1 to 3.
[…]
@@ 5.3 Representation-Level Generalization Analysis
To assess how well the learned latent space separates known and unseen generators, we evaluate the frozen embeddings of Proto-LeakNet’s ResNet18 encoder rather than its classifier head. This isolates the representational geometry induced by signal-leak cues, independent of class supervision. A kernel density estimator, introduced in Section 3, is fitted on the closed embeddings and used to measure how unseen generators align with, or deviate from, the learned manifold under three attention configurations. As shown in Table 2 and visualized in Figure 6, symmetric configurations (Both Off and Both On) yield overlapping score distributions between closed and open samples. Without attention, the encoder loses its ability to emphasize generator-specific latent dimensions, resulting in mixed distributions. When attention is applied to both domains, the same weighting pattern is projected onto unseen data, falsely aligning them with the closed distribution and reducing separability. The asymmetric Dual Attention configuration, where attention is active only for closed embeddings, preserves generator-specific latent biases and produces distinct, non-overlapping densities indicative of true representation-level generalization. All sanity checks, including label shuffling, seed variation, bandwidth cross-validation, and per-channel normalization, confirmed that the observed separation originates from genuine distributional differences rather than data leakage or training artifacts.
Configuration AUC (%) EER OVL
Both Off 57.24 0.44 0.89
Both On 56.62 0.45 0.90
Dual 100.00 0.00 0.00
Figure 6: Plots of both off, both on and dual settings clusters, showing how separation differs from setting to setting.
@@ 5.4 Ablation Study
To assess the contribution of each key component in Proto-LeakNet, we performed three ablation experiments targeting the modules that define the distinct behavior of the pipeline. First, we analyzed the impact of prototypes and feature-attention on latent-space organization and discriminability. As reported in Table 3, removing prototypes substantially degrades the latent structure, reducing both Top-1 Accuracy and Macro AUC, while removing attention diminishes the model’s ability to emphasize the most informative latent dimensions. The combination of both components yields the best overall results, confirming that prototypes and attention jointly enforce structured and discriminative embeddings that better capture signal-leak patterns. We then evaluated the influence of the backbone architecture (Table 4). Replacing the ResNet18 encoder with larger models such as EfficientNet-B4, ViT-B16. ResNet50 or ResNet101 results in lower accuracy and AUC, suggesting that overly deep or transformer-based architectures may overfit local latent variations and weaken generalization. ResNet18 achieves the best balance between representational compactness and discriminative power, providing the most stable feature extraction across post-processed samples. Finally, we investigated the effect of varying the number of prototypes per class ($M$) in Table 5. Using only two prototypes limits intra-class flexibility, while increasing to six introduces redundancy and overlapping prototype regions. Empirically, $M=4$ offered the optimal compromise, yielding compact yet well-separated clusters and the highest closed-set performance. It is important to note that temporal attention pooling, used to aggregate the diffusion-step features, was not ablated. Unlike the feature-attention module within the classifier, the temporal attention pooling provides the only mechanism for combining multi-step latent representations. Removing it would eliminate any temporal aggregation and collapse the entire data flow, making comparison with the original architecture meaningless. Finally, in the first row of Table 3 we reported the results of switching to Stable Diffusion XL instead of Stable Diffusion 2.1; the results reported are almost identical proving that the signal-leak is directly correlated to the forward diffusion process typical of Stable Diffusion. Together, these results validate the architectural choices adopted in Proto-LeakNet and demonstrate how each component contributes to robust and interpretable latent-space attribution.
Table 3: Ablation study on Proto-LeakNet components. We report Top-1 Accuracy and Macro AUC to quantify the effect of removing prototypes and attention mechanisms or changing the Stable Diffusion model.
Experiment Top-1 Acc (%) Macro AUC (%)
SDXL 82.55 98.09
No prototypes 72.63 93.09
No attention 81.23 97.49
No prototypes & no attention 79.80 96.67
Full Proto-LeakNet 82.60 98.13
@@
Backbone Top-1 Acc. (%) Macro AUC (%)
EfficientNet_B4 65.80 91.82
ViT-B16 72.07 94.53
ResNet50 81.83 97.39
ResNet101 76.53 95.36
ResNet18 82.60 98.13
@@
Prototypes ($M$) Top-1 Acc. (%) Macro AUC (%)
$M=2$ 81.17 97.41
$M=6$ 81.33 97.58
$M=4$ 82.60 98.13
6 Discussion
Our study shows that the signal-leak bias is a stable and exploitable forensic cue for source attribution across generative models. Proto-LeakNet is, to our knowledge, the first framework to leverage these latent residuals as discriminative features, shifting the focus from bias correction to exploitation. Across all experiments, the model maintained stable attribution under post-processing and perturbations, confirming that latent-domain methods [latenttracer2024, occlip2024, vasilcoiu2025latte] better preserve generator-specific cues than pixel-space approaches. Despite its robustness, Proto-LeakNet presents several limitations. It requires a distinct evaluation phase for unseen generators and is sensitive to hyperparameter settings such as the loss weighting, attention configuration, and kernel bandwidth in the KDE analysis. Furthermore, the model relies on the assumption that latent signal-leak traces persist across diffusion steps, which may vary across architectures or training regimes. Its prototype–attention head, although interpretable, increases both memory usage and inference time compared to lightweight convolutional baselines. Finally, while the method demonstrates clear separability in representation space, it does not yet provide end-to-end open-set discrimination, which remains an open direction for future research.
@@ 7 Conclusions and Future Works
We introduced Proto-LeakNet, a signal-leak-aware attribution framework that combines robustness and interpretability within a latent-space formulation. By operating directly on diffusion latents and modeling residual generator-specific cues, the model learns a structured embedding space shaped by prototypes and attention, enabling reliable attribution even under strong post-processing. Experiments showed that signal-leak bias is a stable and discriminative forensic cue across both diffusion and non-diffusion models, while density-based evaluation revealed clear separability between known and unseen generators. Beyond accuracy, Proto-LeakNet provides transparent prototype- and feature-level explanations linking latent statistics to interpretable evidence. Future work will explore open-set discriminative training, contrastive regularization for improved disentanglement, and lightweight runtime variants for real-time forensics. Treating signal-leak bias as an interpretable fingerprint rather than a nuisance establishes a foundation for robust and interpretable attribution of generative media.

### v2

@@ 4.1 Evaluation Metrics
    $\text{EER}=\min_{\delta}\big|\text{FAR}(\delta)-\text{FRR}(\delta)\big|,$ (18)
    where $\delta$ denotes the decision threshold. Lower EER indicates better separation between known and unknown samples.
@@
Table 1: AUC (%) results for each generator class on the raw Closed Set. The table reports both per-class AUC and overall Macro AUC, results in bold are the best, underlined ones are second best.
[…]
@@ 5 Experimental Results
We evaluate Proto-LeakNet across multiple scenarios to assess its attribution accuracy, robustness to post-processing, and generalization to unseen data. Our method gets compared against the three best models in WILD [bongini2025wild], FreqNet [freqnet2023], SuSy [susy2024], LatentTracer [latenttracer2024], OCC-CLIP [occlip2024], NPR [npr2024] and LATTE [vasilcoiu2025latte]. Given the similarity in scope and operating domain, we adapted LATTE for source attribution by extending its binary detection objective to a multi-class setting. Owing to its latent-space formulation, the model is naturally predisposed to capture generator-specific cues, making it a suitable baseline for origin attribution. Finally, we include ablation studies to assess the contribution of each component in our pipeline.
@@ 5.1 Results on Closed Set
Proto-LeakNet was first evaluated under the closed-set configuration to verify its ability to learn and capture generator-specific signal-leak patterns. All methods to which we compare were trained in our exact conditions. We first analyze the per-class AUCs reported in Table 1. Although Proto-LeakNet does not achieve the top score for every generator, it consistently ranks among the best across all classes, leading to the highest overall Macro AUC. This indicates stronger balance and generalization in capturing generator-specific signal-leak patterns under the closed-set configuration. Compared to all baselines, Proto-LeakNet attains superior latent discriminability and structured separation. The joint effect of temporal attention and prototype supervision is crucial: attention focuses on the most informative diffusion timesteps, while prototypes enforce geometric consistency by attracting samples toward class centroids and repelling those of different generators. Fig. 3 further shows the Top-1 accuracy distributions, where SuSy [susy2024] slightly surpasses our model on raw images (83.32% vs. 82.60%), yet Proto-LeakNet, due to its highest Macro AUC, yields a more compact and interpretable latent manifold as supported by Fig. 2, effectively encoding generator-specific signal-leak cues that support robust and clear attribution.
Table 2: AUC (%) results for each generator class under increasing post-processing levels (Steps 1–3). The last column reports Macro AUC for all models. Results in bold are the best, underlined ones are second best.
[…]
Configuration AUC (%) EER OVL Attentions Off 57.24 0.44 0.89 Attentions On 56.62 0.45 0.90
@@ 5.2 Results on Post-Processed Closed Set
To assess robustness against real-world degradations, all models and Proto-LeakNet were evaluated on progressively post-processed closed-set samples using pretrained checkpoints from the raw configuration. During this test perturbations from Steps 1-3 (Section 4) were introduced only at inference time, without any fine-tuning. Table 2 reports the per-class AUCs, showing that although Proto-LeakNet does not lead every class, it maintains consistently high performance across all generators compared to the other methods, resulting in the best overall Macro AUC at every degradation level. This confirms stronger resilience to visual distortions compared to image-domain baselines, whose performance progressively collapses with increased perturbation. The robustness of Proto-LeakNet stems from its reliance on latent-domain signal-leak cues, which persist even when pixel-level textures are heavily degraded. Fig. 3 further illustrates the Top-1 accuracies across degradation levels. Methods who are based on latent analysis such as OCC-CLIP, LatentTracer, LATTE and Proto-LeakNet obtain the higher metrics compared to pixel-based methods. In detail, Proto-LeakNet achieves the best results overall, being, 75.89%, 74.51%, and 69.52% for Steps 1, 2, and 3 respectively, while the second best method (LATTE), scores 75.64%, 74.27% and 68.60%. In general, these results demonstrate that Proto-LeakNet preserves generator attribution under aggressive post-processing by capturing stable latent biases rather than superficial image artifacts.
Figure 3: Histogram of the Top-1 accuracy distributions per class for each method, from raw to step 1-3.
@@ 5.3 Representation-Level Generalization Analysis
To assess how well the learned latent space separates known and unseen data, we evaluate the frozen embeddings of Proto-LeakNet’s ResNet18 encoder rather than its classifier head. This isolates the representational geometry induced by signal-leak cues, independent of class supervision. A kernel density estimator (Section 3.4) is fitted on the closed embeddings and used to measure how open-set unseen generators (Section 4) align with, or deviate from, the learned manifold under three attention configurations. As shown in Fig. 4 (a) and Table 3, in the configuration without attention applied in both embeddings the encoder loses its ability to emphasize generator-specific latent dimensions, resulting in mixed distributions. When attention is applied to both domains as illustrated in Fig. 4 (b), the same weighting pattern is projected onto unseen data, falsely aligning them with the closed distribution and reducing separability. The asymmetric configuration (Fig. 4 (c)), where attention is active only for closed embeddings, preserves generator-specific latent biases and produces distinct, non-overlapping densities indicative of true representation-level generalization. To further validate these observations, we employed the CELEBA-HQ dataset [liu2015faceattributes, karras2018progressive], which contains 30,000 high-quality real face images. With the previous experimental configuration (attention applied only on closed embedding), we treated CELEBA-HQ samples as the open-set domain to test whether the learned embeddings would overlap with the closed generators. The resulting latent distributions (Fig. 4 (d)) remained well separated from the closed clusters, confirming that the learned embedding space effectively discriminates unseen, real data without retraining.
Figure 4: Impact of attention configurations on latent-space separation. (a) Disabling attention for both domains leads to overlapping closed and open clusters. (b) Enabling attention for both aligns open embeddings with the closed distribution. (c) The asymmetric setup, attention active only for closed samples, produces well-separated clusters. (d) Applying the same setup as (c) with real CELEBA-HQ images, separability is preserved.
@@ 5.4 Ablation Study
To evaluate the contribution of each component in Proto-LeakNet, we conducted three ablation studies targeting its key modules. First, we examined the role of prototypes and feature-attention in structuring the latent space (Table 4). Removing prototypes substantially disrupts latent organization, lowering both Top-1 Accuracy and Macro AUC, while omitting attention reduces focus on informative latent dimensions. Their combination yields the best results, confirming that prototypes and attention jointly enforce structured, discriminative embeddings that capture signal-leak cues. Replacing Stable Diffusion 2.1 with Stable Diffusion XL (SDXL) produced nearly identical results (Top-1 Acc. 82.55% and Macro AUC of 98.09%), confirming that signal-leak behavior is intrinsic to the diffusion process itself.
Next, we assessed the impact of the encoder architecture (Table 5). Replacing ResNet18 with larger backbones such as EfficientNet-B4, ViT-B16, ResNet50, or ResNet101 decreases performance, indicating that deeper or transformer-based models overfit local latent variations and weaken generalization. ResNet18 provides the best trade-off between compactness and discriminative power, ensuring stable feature extraction across degradations.
We then varied the number of prototypes per class $M$ (Table 6). Using only two prototypes limits intra-class flexibility, while six introduces redundancy and overlap; $M=4$ offers the best balance, yielding compact, well-separated clusters and the highest closed-set performance. Temporal attention pooling, responsible for aggregating diffusion-step features, was not ablated since removing it would collapse temporal aggregation entirely. Moreover, these results validate the design of Proto-LeakNet and highlight how each component contributes to robust and interpretable latent-space attribution.
Table 4: Ablation study on Proto-LeakNet components. We report Top-1 Accuracy and Macro AUC to quantify the effect of removing prototypes and attention mechanisms.
Experiment Top-1 Acc (%) Macro AUC (%) No prototypes 72.63 93.09 No attention 81.23 97.49 No prototypes & no attention 79.80 96.67 Full Proto-LeakNet 82.60 98.13
@@
Backbone Top-1 Acc. (%) Macro AUC (%) EfficientNet_B4 65.80 91.82 ViT-B16 72.07 94.53 ResNet50 81.83 97.39 ResNet101 76.53 95.36 ResNet18 82.60 98.13
@@
Prototypes ($M$) Top-1 Acc. (%) Macro AUC (%) $M=2$ 81.17 97.41 $M=6$ 81.33 97.58 $M=4$ 82.60 98.13
6 Discussion and Limitations
Our study demonstrates that the signal-leak bias constitutes a stable and exploitable forensic cue for source attribution across generative models. Proto-LeakNet represents, to our knowledge, the first framework that explicitly models these latent residuals as discriminative features while providing interpretability across multiple diffusion steps. Across all experiments, the model maintained consistent attribution under post-processing and perturbations, confirming that latent-domain approaches [latenttracer2024, occlip2024, vasilcoiu2025latte] preserve generator-specific cues more effectively than pixel-space methods. Despite its robustness, Proto-LeakNet still presents certain limitations. First, its evaluation is limited to face-centric datasets, leaving open questions about whether signal-leak cues generalize to other visual domains such as objects, scenes, or non-human subjects. Second, the framework relies on the latent representation of Stable Diffusion’s VAE, and it remains unclear whether similar separability would emerge in latents derived from alternative generative families.
@@ 7 Conclusions and Future Works
We introduced Proto-LeakNet, a signal-leak-aware attribution framework that combines robustness and interpretability within a latent-space formulation. By operating directly on diffusion latents and modeling residual generator-specific cues, the model learns a structured embedding space shaped by prototypes and attention, enabling reliable attribution even under strong post-processing. Experiments showed that signal-leak bias is a stable and discriminative forensic cue across both diffusion and non-diffusion models, while density-based evaluation revealed clear separability between known and unseen generators. Beyond accuracy, Proto-LeakNet provides transparent prototype- and feature-level explanations linking latent statistics to interpretable evidence. Future work will extend the evaluation to new visual domains and test alternative generative architectures, such as StyleGAN, to assess whether signal-leak traces persist across different encoder families. Treating signal-leak bias as an interpretable fingerprint, establishes a foundation for robust and interpretable attribution of generative media.